solve by the elimination method
5r-3s=-6
3r+5s=44

Answer by mananth(16949)

5r-3s=-6
3r+5s=44
..
mu;tiply firdt equation by 5 & second equation by 3 and add
5(5r-3s)+3(3r+5s)=-30+132
25r-15s+9r+15s=102
34r=102
divide by 34
34r/34 =102/34
r=3
..
plug the value of r in the first equation
5*3-3s=-6
15-3s=-6
add -15 to both sides
15-15-3s=-15-6
-3s=-21
divide by -3
-3s/-3=-26/-3
s=21/3
s=7